# Evaluating LLMs at Evaluating Temporal Generalization

### v1

@@ Abstract
The rapid advancement of Large Language Models (LLMs) highlights the urgent need for evolving evaluation methodologies that keep pace with improvements in language comprehension and information processing. However, traditional benchmarks, which are often static, fail to capture the continually changing information landscape, leading to a disparity between the perceived and actual effectiveness of LLMs in ever-changing real-world scenarios. Furthermore, these benchmarks do not adequately measure the models’ capabilities over a broader temporal range or their adaptability over time. We examine current LLMs in terms of temporal generalization and bias, revealing that various temporal biases emerge in both language likelihood and prognostic prediction. This serves as a caution for LLM practitioners to pay closer attention to mitigating temporal biases. Also, we propose an evaluation framework Freshbench  for dynamically generating benchmarks from the most recent real-world prognostication prediction. Our code is available at https://github.com/FreedomIntelligence/FreshBench. The dataset will be released soon.
@@ 1 Introduction
Proprietary large language models (LLMs) such as GPT-4 (OpenAI, 2023), Bard (GoogleAI, 2023), Claude (Anthropic, 2023), and PaLM (Anil et al., 2023) has been paralleled by efforts in the open-source community to democratize LLMs. Meanwhile, the rapid push for LLM democratization (Touvron et al., 2023) emphasizes the necessity for dynamic and robust benchmarks that accurately reflecting their evolving capabilities. Traditional benchmarks, broadly classified into knowledge-based assessments and open-dialogue evaluations, each present unique challenges. Knowledge-based assessments, exemplified by MMLU (Hendrycks et al., 2020) and C-Eval (Huang et al., 2023), are prone to data manipulation, raising concerns about their real-world applicability and the risk of data leakage. Open-dialogue evaluations like MT-Bench (Zheng et al., 2023) and Alpaca-Eval (Li et al., 2023), rely on subjective human or model-based judgments, making their outcomes susceptible to the influence of dialogue structure, and thereby, potentially compromising assessment validity.
@@
We argue that the assessment of LLMs should be conducted in scenarios that are resistant to hacking manipulation, with results that are objective and free from evaluator biases (Chen et al., 2024). Additionally, time as a dynamic dimension, can also ensure robust and relevant evaluations. This paper explores an alternative way to evaluate large language models, i.e., on temporal generalization where time is considered as an important aspect. By doing so, the inaccessibility of future content would largely reduce the data contamination issue. Moreover, a retrospective perspective on time could beneficial for temporal bias analysis for LLMs.
@@
To achieve this objective, we explore two specific scenarios designed to measure model capabilities without data leakage and to assess model performance on tasks relevant to both past and future contexts. The first scenario involved generating text from recent sources, such as arxiv papers, news articles, or new books, and Wikipedia. We expected the models to generalize to these diverse, newly generated texts with high language likelihood, thereby supporting the concept of ”intelligence as compression” in a manner free from data leakage. The second scenario focused on predicting future events, suggesting that a superior model should be able to understand the present context, incorporate world knowledge, and generalize to future scenarios. In summary, the evaluation at temporal generalization offers two main benefits: it eliminates the possibility of data leakage through temporal boundaries, and its results can be objectively assessed without reliance on subjective interpretations.
@@
Our contributions are as follows:
  • We thoroughly define and quantify temporal generalization and bias, establishing a basis for understanding and assessing the capabilities of LLMs over time.
  • Through targeted experimental research, we assess the performance of existing large language models (LLMs) across various timeframes. We evaluate how these models handle tasks from different periods, providing insights into their historical and ongoing performance stability and adaptability.
  • We proposed an evaluation framework Freshbench to test the capabilities of LLMs in predicting future events. This framework is dynamically updated to reflect the latest data, ensuring that our evaluations remain relevant and accurate in the face of rapidly evolving data environments.
@@ 2 Conceptualization of Temporal Generalization
In this section, we will conceptualize Temporal Generalization in Sec. 2.1 and further measures it in 2.2. From a retrospective perspective in the Temporal Generalization, we also conceptualize Temporal Biases in Sec. 2.3 and quantify them in 2.4.
@@ 2.1 Temporal Generalization
LLMs face the ongoing challenge of remaining applicability as new text evolves. Consequently, it is crucial for these models to not only understand and generate text based on past and present data but also to anticipate and adapt to future linguistic trends. This leads us to an important concept below.
@@ 2.2 Assessing Temporal Generalization
Having defined temporal generalization, we now turn to a dual-case study to illustrate the dynamic capabilities and limitations of LLMs while adapting to time-evolving linguistic and predictive challenges.
2.2.1 Case I: Language Likelihood
Predicting the likelihood of a sequence of words is a foundational aspect of natural language processing involved in language modeling. Perplexity, a measure of how uncertain a language model is about its predictions, has been critically evaluated in various studies. Saunshi et al. (2021) demonstrates how perplexity can significantly enhance zero-shot inference capabilities, highlighting the importance of these models in understanding and generating language. Wettig et al. (2023) and Tay et al. (2022) have scrutinized the direct correlation between low perplexity during pre-training and improved fine-tuning outcomes, while Liang et al. (2023) discusses the potential issues of data contamination when relying solely on historical text data.
Language Likelihood
In the assessment of language models, metrics such as perplexity (PPL) and bits per character (BPC) offer insights into model performance. Perplexity measures the model’s ability to predict a sequence of words by calculating the exponential of the negative average log likelihood. However, we refine our focus to BPC for a more normalized comparison.
Unlike traditional BPC which is specific to character-level language models, our adapted Bits Per Character (BPC) metric offers a comparison between models by normalizing the log likelihood based on character count, irrespective of the tokenization method, allowing for a more equitable assessment of model performance across different languages and tokenization schemes, the BPC is calculated as follows:
$BPC(T)=\frac{-\sum_{i=1}^{N}\log p(w_{i}|w_{1},\ldots,w_{i-1})}{\text{len-utf-8}(T)}$ (1)
@@ 2.2.2 Case II: Prognostication Prediction
Predictive modeling in the realm of LLMs is increasingly becoming a focal point of research due to its implications for understanding and forecasting dynamic real-world events. Bonde (2022) employed GPT-3 in a few-shot setting to tackle binary questions from the Metaculus platform, demonstrating the capabilities of LLMs in deriving informed predictions. Further, Zou et al. (2022) enriched the discourse by introducing a comprehensive dataset that includes forecasting questions alongside a news corpus. This approach enhances model training by incorporating varied and relevant data contexts. Moreover, Halawi et al. (2024) showcased how integrating news content can significantly improve the accuracy of predictions in real-time scenarios.
Building on this foundation, our research seeks to explore the temporal generalization capabilities of LLMs—how well they adapt and remain accurate as new data emerges and contexts evolve. We aim to deepen the understanding of how these models integrate and apply learned knowledge from past data to predict future events effectively through the case below.
Prognostication Prediction
@@
As an example of using LLMs to predict political outcomes, we gather new factual data about future events, such as "What will President Biden’s approval rating be as of 7 June 2024, according to FiveThirtyEight?". To ensure objectivity in our predictions, answers are typically formatted as ranges within a prediction market scenario. For instance, President Biden’s approval rating might be categorized into intervals such as lower than 34%, at least 34.0%, but less than 36.0%, at least 36.0%, but less than 38.0%, up to 46.0% or higher.
@@
However, since it is impossible to collect actual future data, we adopt a retrospective study approach. We use data available after the event’s question was initially posed and before the prediction market’s closure, treating this as ”future” data for model evaluation. This evaluation is an ongoing, online process, allowing to evaluate the model’s predictive capability in an objective and structured manner. Given the assessment’s nature, we define the accuracy of predictions, denoted by $Acc$, as follows:
$Acc=\frac{N_{\text{correct}}}{N_{\text{total}}}$ (2)
where $N_{\text{correct}}$ represents the number of correct predictions and $N_{\text{total}}$ is the total number of predictions made. This metric evaluates the effectiveness of LLMs in predicting future events, serving as a straightforward measure of their performance in real-world scenarios.
This approach allows us to continuously evaluate the model with the latest data, ensuring that the assessment is reflective of the model’s current capabilities and the evolving nature of the information landscape. By evaluating the generalization capabilities of LLMs across those scenarios, we seek to better understand how these models incorporate and reflect perceived world knowledge when confronted with dynamic, real-world conditions.
@@ 2.3 Temporal Bias
In the context of LLMs, there exists two biases, namely Nostalgia Bias (see Definition 2.2) and Neophilia Bias (see Definition 2.3), which reflect two contrasting tendencies in how these models might predict or generate text about the past, present, or future. Understanding the two biases is crucial for interpreting the models’ outputs, especially when attempting to predict future text. A visual representation of these biases is provided in Figure 1, where the trends associated with nostalgia and neophilia biases are plotted over time.
Figure 1: Schematic figure of nostalgia bias and neophilia bias. Prior to the cutoff time, the performance of the model exhibiting nostalgia bias is skewed toward past dates, showing a preference for historical data. In contrast, a model with neophilia bias is skewed toward future dates, indicating a tendency to favor more recent information. However, after the cutoff time, the model’s understanding of future trends declines, leading to a decrease in performance.
Definition 2.2.
Nostalgia Bias Nostalgia Bias in LLMs refers to a tendency to over-rely on or give undue preference to historical data or past events in generating text. This means that when predicting text about the past or the present, the model might disproportionately reflect the perspectives, styles, or information prevalent in its training data, which stops at a certain point in the past. This bias can lead to outputs that are more reflective of past conditions and less attuned to recent developments or changes that have occurred since the model’s last training data update. In essence, Nostalgia Bias can make a model seem ‘stuck in the past’, limiting its ability to accurately represent the present or predict the future.
Definition 2.3.
Neophilia Bias Neophilia Bias, on the other hand, would imply a model’s overemphasis on novelty, future trends, or speculative ideas, potentially at the expense of well-established information from the past or present. This bias could manifest in a model’s predictions or text generation leaning towards overly optimistic or speculative futures without sufficient grounding in historical data or current realities. While less likely in practice due to the nature of LLM training on historical datasets, a perceived Neophilia Bias could arise from the way prompts are structured or the interpretation of the model’s outputs, especially if they seem to prioritize futuristic concepts over historical accuracy.
In a word, Nostalgia bias refers to the model’s tendency to reflect knowledge from earlier periods before the training cutoff, whereas neophilia bias indicates a model’s inclination towards the most recent information near the cutoff time.
Challenges in Temporal Generalization Predicting future text with LLMs while navigating Nostalgia Bias and Neophilia Bias presents distinct challenges. On one hand, Nostalgia Bias can cause models to generate predictions that might not fully account for recent trends or emerging phenomena, making the predictions potentially outdated or irrelevant. On the other hand, an overcorrection towards Neophilia Bias (or an attempt to counteract Nostalgia Bias) could lead to predictions that are too speculative or disconnected from past (especially long-term) trajectories. The challenge lies in balancing these biases to leverage the historical and contemporary data effectively, ensuring that predictions about the future are both informed by past patterns and open to new possibilities. Achieving this balance requires careful selection and adjustment of the model’s training data, and possibly the integration of up-to-date information alongside the model’s outputs.
2.4 Qualitative Measurement of Temporal Bias
We aim to assess temporal bias qualitatively by introducing the Temporal Bias Index (TBI).
Definition 2.4.
Temporal Bias Index (TBI) qualitatively determines whether the model’s performance is biased towards past or future dates. It could be calculated by analyzing the slope of the performance over time, which indicates changes in model performance. A negative slope suggests a decline in performance, indicating a bias towards historical data, whereas a positive slope indicates an improvement, suggesting a bias towards more recent data.
The methodology for calculating TBI varies based on the metric used to measure model performance. Specifically, when performance is measured by language likelihood, we calculate the slope of BPC over time directly: This slope indicates whether BPC is increasing (a positive slope, suggesting degradation) or decreasing (a negative slope, indicating improvement).
Conversely, when performance is measured by the correctness of answers to questions, we define several key symbols: $n$ represents the total number of attempts at questions within the month, $P$ is the observed frequency of correct answers, $E(P_{\text{random}})$ is the expected frequency of correct answers under random guessing calculated based on the average number of options, and $\Phi$ is the cumulative distribution function (CDF) of the standard normal distribution.
Using these symbols, the performance evaluation formula TBI is presented as follows:
$\text{TBI}=\Phi\left(\frac{\sqrt{n}\times(P-E(P_{\text{random}}))}{\sqrt{P\times(1-P)}}\right)$ (3)
Equation 3 normalizes the deviation between the observed frequency of correct answers and the expected frequency under random guessing, producing a performance score between 0 and 1. A higher score indicates better performance relative to the expected level of a standard normal distribution, effectively adjusting for the biases introduced by random guessing and the variability in the number of options.
Figure 2: A case study of temporal bias, assessing the accuracy of responses to queries from Good Judgment Open dataset (smoothed), spanning from September 1, 2015, to the models’ cutoff time. The blue line for ”Yi-6B” is right-skewed with a bias of -0.006, signifying a tendency to favor past dates, which infers a ’Nostalgia’ bias within the model. In contrast, the green line, which depicts the ”Phi-1.5” model, shows a minimal bias of 0.001, signifying a slight preference for more recent information. This can be interpreted as a ’Neophilia’ bias.
In our case study of temporal bias, as depicted in Figure 2, we demonstrate the validity of our defined measurement of temporal bias. The bias metric effectively gauges whether the model’s performance exhibits a tendency to favor either past or future dates.
3 Temporal Generalization in terms of Language Likelihood
3.1 Experimental Settings
@@
Our system aggregates data from various platforms, including Yahoo for financial news, BBC for political insights, Reddit for discussions, Wattpad for literature, Wikipedia for encyclopedic updates, arXiv for academic research, GitHub for software trends, and Quora for diverse questions and answers. Utilizing a Python-based crawling framework, detailed in our public repository, it’s designed for adaptability and efficient data collection, with Playwright addressing navigation and specific site challenges. We also did some basic data pre-processing, see details in Appendix A. An overview of the data volume and the average entry length from a single crawl is provided in Appendix B.
@@
To further delve into the model’s performance dynamics over varying text lengths, we truncate the input text to the first 5000 characters and compute the sum of loss. The loss values are then accumulated to derive their sum, providing insight into the model’s efficiency at different stages of text comprehension.
Additionally, we calculate the average loss over token length milestones—specifically at 10, 50, 100, 500, 1000, 1500, and 2000 tokens. This step allows us to observe the impact of length on the model’s understanding and predictive capabilities, highlighting how performance metrics vary with increasing input size.
3.2 Temporal Bias
The slope is obtained by performing a linear regression analysis where the independent variable $x$ is the month and the dependent variable $y$ is the model’s performance score, here is BPC, the lower the better, a negative slope indicates the presence of Neophilia Bias. Conversely, a positive slope would suggest that the model’s performance is decreasing, which shows Nostalgia Bias.
To delve deeper into the temporal generalization capabilities of models, we plot the BPC curve of fresh text over time. We sampled arXiv papers from 2015 to 2024 and applied smoothing techniques to the data because the number of samples was not sufficient to completely stabilize the average language likelihood. Nonetheless, a slight downward trend in model BPC can be observed, which is noteworthy. What is more, models with initially higher BPC showed a larger percentage decrease in BPC (as seen in models like OPT-2.7B and PHI-2 in Figure  4). For models with lower initial BPC, the decrease trend is less pronounced. Taking arXiv text as an example, nearly most models do not show clear flaws in absolute generalization, as BPC decreases over time. Moreover, in terms of arXiv data, LLMs exhibit a neophilia bias rather than a nostalgia bias, capturing the present more effectively than the past.
Turning to BBC news text, the observed trends in BPC from 2020 to 2023, depicted in Figure 4, provide insight into how well these models process and understand news content over a span of several years. As delineated in the figure, we note a pronounced increase in BPC across the observed period, which suggests a diminishing adaptability of the models to newer data, indicative of a nostalgia bias. This trend may stem from the dynamic and evolving nature of news narratives. Detailed data for arXiv, BBC, and Wiki datasets are listed in Table 1.

### v2

# Is Your LLM Outdated? Evaluating LLMs at Temporal Generalization

@@ Abstract
The rapid advancement of Large Language Models (LLMs) highlights the urgent need for evolving evaluation methodologies that keep pace with improvements in language comprehension and information processing. However, traditional benchmarks, which are often static, fail to capture the continually changing information landscape, leading to a disparity between the perceived and actual effectiveness of LLMs in ever-changing real-world scenarios. Our study examines temporal generalization, which includes the ability to understand, predict, and generate text relevant to past, present, and future contexts, revealing significant temporal biases in LLMs. We propose an evaluation framework, for dynamically generating benchmarks from recent real-world predictions. Experiments demonstrate that LLMs struggle with temporal generalization, showing performance decline over time. These findings highlight the necessity for improved training and updating processes to enhance adaptability and reduce biases. Our code, dataset and benchmark are available at https://github.com/FreedomIntelligence/FreshBench.
@@ 1 Introduction
Proprietary large language models (LLMs) such as GPT-4 [1], Bard [2], Claude [3], and PaLM [4] has been paralleled by efforts in the open-source community to democratize LLMs. Meanwhile, the rapid push for LLM democratization [5] emphasizes the necessity for dynamic and robust benchmarks that accurately reflect their evolving capabilities. Traditional benchmarks, broadly classified into knowledge-based assessments and open-dialogue evaluations, each present unique challenges. Knowledge-based assessments, exemplified by MMLU [6] and C-Eval [7], are prone to data manipulation, raising concerns about their real-world applicability and the risk of data leakage. Open-dialogue evaluations like MT-Bench [8] and Alpaca-Eval [9], rely on subjective human or model-based judgments, making their outcomes susceptible to the influence of dialogue structure, and thereby, potentially compromising assessment validity.
Given the limitations of traditional methods, researchers are exploring new ways to measure model performance more comprehensively. Recently, several methods have been proposed to evaluate models at different times and in various contexts[10, 11, 12]. TimeQA[13] evaluates a model’s ability to handle temporal evolution, while Situated QA[14] focuses on a model’s performance in specific contexts. StreamingQA[15] introduces the concept of continuous information flow, while RealTimeQA[11] and FreshQA[12] emphasize real-time and up-to-date information processing capabilities. However, existing methods may be prone to data leakage during their evaluation processes and typically focus on current or past data, lacking exploration of future event prediction capabilities.
@@
We believe that the evaluation of LLMs should be conducted in scenarios that are resistant to hacking manipulation, with results that are objective and free from evaluator biases [16]. Additionally, time as a dynamic dimension, can also ensure robust and relevant evaluations. This paper explores an alternative way to evaluate large language models, i.e., on temporal generalization where time is considered as an important aspect. By doing so, the inaccessibility of future content would largely reduce the data contamination issue. Moreover, a retrospective perspective on time could be eneficial for temporal bias analysis for LLMs.
@@
To achieve this objective, we explore two specific scenarios designed to measure model capabilities without data leakage and to assess model performance on tasks relevant to both past and future contexts. The first scenario involved generating text from recent sources, such as arxiv papers, news articles, or new books, and Wikipedia. We expected the models to generalize to these diverse, newly generated texts with high language likelihood, thereby supporting the concept of "intelligence as compression" in a manner free from data leakage. The second scenario focused on predicting future events, suggesting that a superior model should be able to understand the present context, incorporate world knowledge, and generalize to future scenarios. In summary, the evaluation at temporal generalization offers two main benefits: it eliminates the possibility of data leakage through temporal boundaries, and its results can be objectively assessed without reliance on subjective interpretations.
@@
Our contributions are as follows: 1) We thoroughly define and quantify temporal generalization and bias, establishing a basis for understanding and assessing the capabilities of LLMs over time. 2) We proposed an benchmark Freshbench to test the capabilities of LLMs in predicting future events. This benchmark is dynamically updated to reflect the latest data, ensuring that our evaluations remain relevant and accurate in the face of rapidly evolving data environments. It experimentally assesses performance of existing LLMs across various timeframes, providing insights into their historical and ongoing performance stability and adaptability.
@@ 2 Conceptualization of Temporal Generalization
Figure 1: Schematic figure of temporal generalization.
In this section, we will conceptualize Temporal Generalization in Sec. 2.1 and further measure it in 2.2. From a retrospective perspective in the Temporal Generalization, we also conceptualize Temporal Biases in Sec. 2.3. In general, prior to the cutoff time, a LLM might exhibit a Nostalgia bias that is skewed toward past dates (i.e., aligning better with historical data) or a Neophilia bias that is skewed toward future dates ( favoring more recent information). However, after the cutoff time, the model’s understanding of future trends declines, leading to temporal degeneration.
@@ 2.1 Temporal Generalization
LLMs face the ongoing challenge of remaining applicability as new text evolves. Consequently, it is crucial for these models to not only understand and generate text based on seen (e.g., past) data but also to anticipate and adapt to future data. This inspires temporal generalization as below.
@@ 2.2 Assessing Temporal Generalization
Based on definition in Sec. 2.1, we turn to assess temporal generalization through the lens of 1) I: compression intelligence in Sec. 2.2.1 or II): prognostication prediction in Sec. 2.2.2.
2.2.1 Case I: Compression Intelligence over Time
In the assessment of compression intelligence, metrics such as perplexity (PPL) and bits per character (BPC) offer insights into model performance. We refine our focus to BPC for a more normalized comparison. Unlike traditional BPC which is specific to character-level language models, our adapted Bits Per Character (BPC) metric offers a comparison between models by normalizing the log likelihood based on character count, irrespective of the tokenization method, allowing for a more equitable assessment of model performance across different languages and tokenization schemes, the BPC is calculated as follows:
$\small BPC(T)=\frac{-\sum_{i=1}^{N}\log p(w_{i}|w_{1},\ldots,w_{i-1})}{\text{len-utf-8}(T)}$ (1)
@@ 2.2.2 Case II: Prognostication Prediction
LLMs could performs prognostication prediction, as seen in [18] which employed GPT-3 in a few-shot setting to tackle binary questions from the Metaculus platform. Further, [19] enriched the discourse by introducing a comprehensive dataset that includes forecasting questions alongside a news corpus. This approach enhances model training by incorporating varied and relevant data contexts. Moreover, [20] showcased how integrating news content can significantly improve the accuracy of predictions in real-time scenarios.
@@
Building on this foundation, our research seeks to explore the temporal generalization capabilities of LLMs: how well they adapt and remain accurate as new data emerges and contexts evolve?. As an example of using LLMs to predict political outcomes, we gather new factual data about future events, such as "What will President Biden’s approval rating be as of 7 June 2024, according to FiveThirtyEight?". To ensure objectivity in our predictions, answers are typically formatted as ranges within a prediction market scenario. For instance, President Biden’s approval rating might be categorized into intervals such as lower than 34%, at least 34.0%, but less than 36.0%, at least 36.0%, but less than 38.0%, up to 46.0% or higher.
@@
However, since collecting actual future data is impossible, we adopt a retrospective study approach. We use data available after the event’s question was posed and before the prediction market’s closure, treating this as "future" data for model evaluation. This ongoing, online evaluation objectively assesses the model’s predictive capability. We define prediction accuracy, denoted by $Acc$, as follows:
$\small Acc=\frac{N_{\text{correct}}}{N_{\text{total}}}$ (2)
where $N_{\text{correct}}$ is the number of correct predictions among $N_{\text{total}}$ future prediction questions. See. Sec. 4.1 for the collected future prediction question set over time. This metric evaluates the effectiveness of LLMs in predicting future events. This approach also allows continuous evaluation with the latest data, ensuring the assessment reflects the model’s current capabilities and the evolving information landscape.
2.3 Temporal Bias in a Retrospective View
Temporal biases are multifaceted in previous literature. For instance, temporal bias can refer to the assumption of an incorrect sequence of events that misleads our reasoning about causality [21] or user preferences over time in point-of-interest recommendations [22]. Recently, biases have raised more challenges in pre-trained language models [23], particularly focusing on under-represented societies. [24] examines lookahead bias in pre-trained language models, providing insights into how these biases affect model performance. In the context of LLMs, temporal bias refers to the preferences of LLMs over time.
There exists two biases, namely Nostalgia Bias (see Definition 2.2) and Neophilia Bias (see Definition 2.3), which reflect two contrasting tendencies in how these models might predict or generate text about the past, present, or future. A visual representation of these biases is provided in Fig. 1, where the trends associated with Nostalgia and Neophilia biases are plotted over time.
Definition 2.2.
Nostalgia Bias refers to the over-reliance on historical data when generating text. This bias causes models to reflect perspectives, styles, or information prevalent in past training data, leading to outputs that may not account for recent developments. Essentially, Nostalgia Bias can make a model seem ‘stuck in the past’, limiting its ability to accurately represent the present or predict the future.
Definition 2.3.
Neophilia Bias implies a model’s overemphasis on novelty and future trends, potentially at the expense of established information. This bias could result in predictions that are overly speculative and not well-grounded in historical data or current realities. Though less common due to LLM training on historical datasets, Neophilia Bias could be perceived from prompt structuring or interpretation of outputs that prioritize futuristic concepts over historical accuracy.
Challenges in Temporal Generalization Predicting future text with LLMs while navigating Nostalgia Bias and Neophilia Bias presents distinct challenges. Nostalgia Bias can lead to outdated predictions, while overcorrection towards Neophilia Bias might result in speculative predictions disconnected from past trajectories. Balancing these biases is crucial to leverage historical and contemporary data effectively, ensuring future predictions are informed by past patterns and open to new possibilities. This requires careful adjustment of the model’s training data and integration of up-to-date information.
3 Temporal Generalization through the Lens of Compression Intelligence
@@ 3.1 Temporal Bias
3.1.1 Protocol for Quantitative Measurement
@@
Our system aggregates data from various platforms, including Yahoo for financial news, BBC for political insights, Reddit for discussions, Wattpad for literature, Wikipedia for encyclopedic updates, arXiv for academic research, GitHub for software trends, and Quora for diverse questions and answers. Utilizing a Python-based crawling framework, detailed in our public repository, it’s designed for adaptability and efficient data collection, with Playwright addressing navigation and specific site challenges. We also did some basic data pre-processing, due to space limitation, details in Appendix B. An overview of the data volume and the average entry length from a single crawl is provided in Appendix C.
@@
Experimental Settings To further delve into the model’s performance dynamics over varying text lengths, we truncate the input text to the first 5000 characters and compute the sum of loss. The loss values are then accumulated to derive their sum, providing insight into the model’s efficiency at different stages of text comprehension. Throughout this experiment, approximately 300 GPU hours on NVIDIA A800 were consumed.
3.1.2 Observation of Temporal Bias in Compression Intelligence
To comprehensively explore the temporal generalization capabilities of models, we plot the BPC curve of fresh text over time. In text data before release, a negative slope indicates the presence of Neophilia Bias. Conversely, a positive slope would suggest that the model’s performance is decreasing, which shows Nostalgia Bias.
We sampled arXiv papers from 2015 to 2024 and applied smoothing techniques to the data because the number of samples was not sufficient to completely stabilize the average language likelihood. Nonetheless, a slight downward trend in model BPC can be observed, which is noteworthy. What is more, models with initially higher BPC showed a larger percentage decrease in BPC (as seen in models like OPT-2.7B and PHI-2 in Figure  3). For models with lower initial BPC, the decrease trend is less pronounced. Taking arXiv text as an example, BPC decreases over time, LLMs exhibit a Neophilia bias rather than a Nostalgia bias, capturing the present more effectively than the past.
Turning to BBC news text, the observed trends in BPC from 2020 to 2023, depicted in Fig. 3, provide insight into how well these models process and understand news content over a span of several years. We note a pronounced increase in BPC across the observed period, which suggests a diminishing adaptability of the models to newer data, indicative of a Nostalgia bias. This trend may stem from the dynamic and evolving nature of news narratives. Detailed result for various sources are in E